OpenStudy (anonymous):

for the polynomial, one zero is given. find all others. p(x)=x^3+2x^2-6x+8; 1+i

OpenStudy (anonymous):

ooh kay now we have a question one zero is \(1+i\) the other is its conjugate \(1-i\) lets find the quadratic with those zeros there are three ways one is annoying one is easy one is really really easy

OpenStudy (anonymous):

I would prefer the very easy way :)

OpenStudy (anonymous):

lets do the easy way work backwards \[x=1+i\] subtract 1\[x-1=i\] square both sides (carefully) \[(x-1)^2=i^2\\ x^2-2x+1=-1\] add 1 \[x^2-2x+2\] is your quadratic

OpenStudy (anonymous):

ok very very easy way is very very easy but it requires memorizing something if the zeros are \(a+bi\) then the quadratic is \[x^2-2ax+a^2+b^2\]

OpenStudy (anonymous):

in your case \(1+i\) you have \(a=1,b=1\) quadratic is \[x^2-2\times 1+1^2+1^2=x^2-2x+2\]

OpenStudy (anonymous):

this means \[p(x)=x^3+2x^2-6x+8=(x^2-2x+2)(something)\] and the "something" should be real easy to find

OpenStudy (triciaal):

@satellite73 thanks for sharing the "very very easy way" don't remember if I learnt that or proves that it's not easy if you don't remember

OpenStudy (anonymous):

yw it is usually the case that the more you do something the easier it becomes, but memorizing something always risks memorizing it incorrectly the very hard way is usually what is taught, multiplying \[(x-(a+bi))(x-(a-bi))\] which is a pain

OpenStudy (triciaal):

oh i know how to do it just don't recall the formula you provided

OpenStudy (triciaal):

i learnt the hard way and yes it comes easy with practice

OpenStudy (anonymous):

@Alexy did you find the part after \[x^3+2x^2-6x+8=(x^2-2x+2)(something)\]?